OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x + y = 11 -3x + 2y = 1

jimthompson5910 (jim_thompson5910):

notice we have 3x in the top equation and -3x in the bottom equation

jimthompson5910 (jim_thompson5910):

they add to 3x + (-3x) = 0x = 0 so the x terms go away when you add 3x and -3x

jimthompson5910 (jim_thompson5910):

add the y terms: y+2y = 3y add the constants on the right side: 11+1 = 12

jimthompson5910 (jim_thompson5910):

if we add the equations, we get 3x + y = 11 -3x + 2y = 1 ------------- 0x + 3y = 12 giving us 3y = 12 Making sense so far?

OpenStudy (anonymous):

Yes, I see now, i accidently made the 12 a 10 i see now

jimthompson5910 (jim_thompson5910):

If 3y = 12 then y = ??